oliviac1111:

What is the midpoint of the x-intercepts of f(x) = (x – 4)(x + 4)? (0,0) (0,4) (–4,0) (2,0)

Florisalreadytaken:

hmm ok, so this is what we have: \[ 0= (x - 4)(x + 4) \] that leads us to 2 different solutions: \( 1) \ \ x-4=0 \Rightarrow x=0+4 \Rightarrow x=4 \) \( 2) \ \ x+4=0 \Rightarrow x=0-4 \Rightarrow x=-4\) what do we do next?

oliviac1111:

oh shoot uhm

oliviac1111:

its asking me where they meet on the graph right ?

Florisalreadytaken:

the MIDDLE where they meet, yes so we have got \[ (4,0) ~and~ (-4,0) \] you can directly tell where the 2 points middle is right?

oliviac1111:

(4,-4)

oliviac1111:

wait

oliviac1111:

-4

Florisalreadytaken:

Nononono they meet in the middle

1 attachment
oliviac1111:

that would just be zero tho

Florisalreadytaken:

exactly -- \( (0,0) \)

oliviac1111:

ohhhh oml lol thank u

oliviac1111:

can u help me with more?

oliviac1111:

@Florisalreadytaken The equation x2 – 1x – 90 = 0 has solutions {a, b}. What is a + b? –19 –9 1 10